# Continuous time asymptotic representations for adaptive experiments

### v1

@@ E.2. Locally REGROW optimal e-processes
As a leading example—following grunwald2024safe—consider the case where each $h_{a}$ is scalar and
@@
for some constant $K<\infty$. In this setting, the optimization problem (E.2) factorizes across arms, yielding a product form solution $w^{*}=\prod_{a}w_{a}^{*}(h^{(a)})$, and the overall channel capacity becomes the sum of individual channel capacities for each arm. Classical results from smith1971 establish that the least-favorable distribution $w_{a}^{*}(\cdot)$ is always discrete. For low signal to noise ratios, specifically, when $K/I_{a}^{1/2}q_{a}\lessapprox 1.4$, $w_{a}^{*}(\cdot)$ reduces to a symmetric two-point prior supported on $\{-K,K\}$. However, when $q_{a}$ falls below the critical threshold $K/(1.4I_{a}^{1/2})$, the support of $w_{a}^{*}$ expands to include more than two points and must be computed via a finite-dimensional convex program, as described in smith1971.

### v2

@@ E.2. Locally REGROW optimal e-processes
As a leading example—following Grünwald et al. (2024)—consider the case where each $h_{a}$ is scalar and
@@
for some constant $K<\infty$. In this setting, the optimization problem (E.2) factorizes across arms, yielding a product form solution $w^{*}=\prod_{a}w_{a}^{*}(h^{(a)})$, and the overall channel capacity becomes the sum of individual channel capacities for each arm. Classical results from Smith (1971) establish that the least-favorable distribution $w_{a}^{*}(\cdot)$ is always discrete. For low signal to noise ratios, specifically, when $K/I_{a}^{1/2}q_{a}\lessapprox 1.4$, $w_{a}^{*}(\cdot)$ reduces to a symmetric two-point prior supported on $\{-K,K\}$. However, when $q_{a}$ falls below the critical threshold $K/(1.4I_{a}^{1/2})$, the support of $w_{a}^{*}$ expands to include more than two points and must be computed via a finite-dimensional convex program, as described in Smith (1971).